Comprehensive Problem 4 a-g Kristin Malone opened Kristins Maids Cleaning Service on July 1, 2019. During July, the company(b) (c) Prepare a trial balance at July 31 on a worksheet. Enter the following adjustments on the worksheet, and complete theKRISTINS MAIDS CLEANING SERVICE Prepare a worksheet. | For the Month Ended July 31, 2019 V Adjustments Adjusted Trial Balanc
Answer #1

1) The required worksheet is shown as follows:- (Amounts in $)

KRISTIN'S MAIDS CLEANING SERVICE
Worksheet
Account Titles Trial Balance Adjustments Adjusted Trial Balance Income Statement Balance Sheet
Dr. Cr. Dr. Cr. Dr. Cr. Dr. Cr. Dr. Cr.
Cash 7,992 7,992 7,992
Accounts Receivable 3,800 1,500 5,300 5,300
Supplies 800 600 200 200
Prepaid Insurance 1,608 134 1,474 1,474
Equipment 11,000 11,000 11,000
Accounts Payable 7,400 7,400 7,400
Common Stock 15,000 15,000 15,000
Dividends 700 700 700
Service Revenue 5,400 1,500 6,900 6,900
Gasoline Expense 300 300 300
Salaries and Wages Expense 1,600 400 2,000 2,000
Totals 27,800 27,800
Depreciation Expense 220 220 220
Accum. Depr.-Equipment 220 220 220
Insurance Expense 134 134 134
Supplies Expense 600 600 600
Salaries and Wages Payable 400 400 400
Totals 2,854 2,854 29,920 29,920 3,254 6,900 26,666 23,020
Net Income 3,646 3,646
Totals 6,900 6,900 26,666 26,666

Working Notes:-

1) Calculation of Cash Balance in Trial Balance (Amounts in $)

Issue of Common Stock 15,000
Cash paid for Truck (3,200)
Prepaid Insurance paid (1,608)
Amount owed for truck paid (1,000)
Amount owed for supplies paid (200)
Salaries paid (1,600)
Collection from Customers 1,600
Gasoline Expense paid (300)
Dividends paid (700)
Ending Balance of Cash before Adjustments 7,992

Accounts Receivable before adjustments = 3,900-1,600+1,500 = $3,800

Accounts Payable before adjustments = (11,000-3,200)+800-1,000-200 = $7,400